OpenStudy (anonymous):

Find the area of a sector with an arc length of 60 in. and a radius of 15 in. A. 117.75 in.² B. 282.60 in.² C. 450 in.² D. 6,750 in.²

OpenStudy (dumbcow):

\[Area = \pi r^2 * \frac{\theta}{2\pi} \] \[Arc length = r \theta \rightarrow \theta = \frac{60}{15} = 4\] \[\rightarrow Area = \pi (15)^2 * \frac{4}{2 \pi} = 450\]